# JWST’s First View of Tidal Disruption Events: Compact, Accretion-Driven Emission Lines & Strong Silicate Emission in an Infrared-selected Sample

### v1

@@ Abstract
Mid-infrared (MIR) emission from tidal disruption events (TDEs) is a powerful probe of the circumnuclear environment around dormant supermassive black holes. This emission arises from the reprocessing of intrinsic emission into thermal MIR emission by circumnuclear dust. While the majority of optical- and X-ray-selected TDEs show only weak dust echoes consistent with primarily unobscured sight lines, there have been growing efforts aimed at finding TDEs in obscured environments using MIR selection methods. In this work, we present the first JWST observations of 4 MIR-selected TDEs with JWST Mid-Infrared Instrument (MIRI) Medium-Resolution Spectrometer (MRS). Two of these sources show flares in other wavelength bands (one in optical, one in X-ray), while the other two are MIR-only transients. None of these TDEs showed pre-outburst nuclear activity, but all of the MIRI/MRS observations reveal emission lines associated with highly ionized gas, implying ionization from TDE accretion. Additionally, all four sources show silicate emission features around 10 and 18 µm  that are much stronger than the features seen in active galactic nuclei (AGN). We suggest that the emitting dust is optically thin to its own emission and show that the MIR spectrum is consistent with emission from optically thin dust along an isodelay surface. All four sources show an excess at short wavelengths ($\lambda<8$ µm), which could arise from a late-time plateau in the intrinsic flare, akin to what is seen in late-time UV observations of unobscured TDEs, although self-consistent dust modeling is required to fully assess the strength of this late-time plateau.
@@ 1 Introduction
In the past 50 years, tidal disruption events (TDEs) have gone from a theoretical concept (Rees, 1988; Evans & Kochanek, 1989; Cannizzo et al., 1990) to an observable phenomenon (for a recent review, see Gezari, 2021), becoming one of the leading methods to study the dominant fraction of dormant supermassive black holes (SMBHs). TDEs occur when a star on a nearly radial orbit gets so close to an SMBH that the tidal forces overcome the star’s self-gravity. A fraction of the star’s mass falls back onto the SMBH, producing a flare of accretion that can be observed across the electromagnetic spectrum. In the late 1990s, the first TDEs were detected with transient soft, thermal X-ray emission, indicative of the formation of an accretion flow (e.g. Bade et al., 1996; Komossa & Bade, 1999; Komossa & Greiner, 1999; Donley et al., 2002; Esquej et al., 2007). A few more TDEs were discovered in the early 2000s with time-domain searches in the ultraviolet (UV) band using GALEX (e.g. Gezari et al., 2006, 2008, 2009). The discovery space is now dominated by time-domain optical surveys, which have increased the discovery rate by a factor of 10 already (van Velzen et al., 2011; Gezari et al., 2012; van Velzen et al., 2021b; Hammerstein et al., 2023; Yao et al., 2023) and are expected to increase the rate again by at least a factor of 10 with next generation surveys like the Legacy Survey for Space and Time (LSST) at the Vera Rubin Observatory (Bricman & Gomboc, 2020).
These canonical detection methods tend to find TDEs in relatively gas- and dust-poor environments, as evidenced by the low covering factors measured with TDE dust echoes in the mid-infrared (MIR; van Velzen et al., 2016; Jiang et al., 2021a). However, in gas- and dust-rich environments, the nuclear dust may have a larger covering factor, making it more likely that the dust covers the sight line to the nucleus. Dust around accreting SMBHs reprocesses the higher energy optical/UV/X-ray photons into thermal MIR emission (van Velzen et al., 2021a), and thus, the MIR may be the only way to detect TDEs in gas- and dust-rich environments. Initial searches for MIR flares in ultraluminous infrared galaxies (U/LIRGs) revealed a handful of TDE-like events (e.g. Mattila et al., 2018; Kool et al., 2020), suggesting an elevated rate of TDEs in obscured environments (Reynolds et al., 2022), although many of these showed some level of activity from an active galactic nucleus (AGN) prior to the MIR flare. Other efforts have involved searching for MIR flares in integrated photometry from the NEOWISE mission (Mainzer et al., 2014), using spectroscopically confirmed galaxies from SDSS (Jiang et al., 2021b). This approach revealed many potential TDE candidates, but was also agnostic to past AGN activity, thereby introducing many AGN contaminants into the sample.
@@ 4.2 Time-Dependent Optically Thin Dust Modeling
where $\overline{Q}_{\nu}$ is the average absorption efficiency integrated over the source spectrum. In the UV and optical band, $\overline{Q}_{\nu}\approx 1$, while $\overline{Q}_{\nu}$ is roughly 2-3 orders of magnitude lower in the IR. For relatively conservative choices of the density profile (e.g. the highest density and $n_{d}(r)\propto r^{-1}$), the dust is indeed optically thin to its own emission for the size scales we are investigating here ($r\sim$ few pc), regardless of the order of magnitude uncertainties on the gas density in galactic centers.
Motivated by the silicate emission features and this order of magnitude estimate, we modeled the JWST MIRI spectra with optically thin dust, neglecting the effects of scattering but accounting for the time-dependent effects from TDEs. For this initial investigation, we adopt dust composed of 53% silicate and 47% graphite grains (i.e. an MRN mixture, typical for the Milky Way’s ISM; Mathis et al., 1977) with a single size of $a=0.1$ µm. This choice of an MRN dust mixture is motivated by the presence of both silicates and PAHs (i.e. carbon-rich dust) in the MIR spectra. We use the astronomical silicate and graphite optical properties from Draine & Lee (1984) and Laor & Draine (1993) [^2] See https://www.astro.princeton.edu/~draine/dust/dust.diel.html. To simplify the initial computation, we assume a $\delta$-function input flare at $t=0$ and compute the resulting isodelay surface seen by the observer at time $t$, given by $t=\frac{r}{c}\left(1-\cos\theta\right)$. Along each point on this surface, we numerically compute the expected temperature of the dust by solving the energy balance equation, given by
@@
where $L$ is the bolometric luminosity emitted by the source at a time $t^{\prime}=t-\frac{r}{c}\left(1-\cos\theta\right)$, $\tau$ is the integrated optical depth given by Equation 2, $Q_{\nu}$ is the absorption efficiency, and $T_{d}(r,\theta,t)$ is the dust temperature. This prescription considers only the radiative cooling of the dust, ignoring cooling from grain sublimation, which is only dominant at significantly higher temperatures than probed here ($T\gtrsim 2800$ K; e.g. Waxman & Draine 2000; Lu et al. 2016). Given the unknown source spectrum and the high efficiency of dust absorption at optical and UV wavelengths, we assume $\overline{Q}_{\nu}=1$. We adopt a simple power-law density profile given by $n(r)=n_{\mathrm{sub}}(r/r_{\mathrm{sub}})^{-\gamma}$, where the subscript “sub” denotes values at the sublimation radius.
Finally, we sum up the resulting IR spectrum over all radii following Dwek (1983) Equation (5), given by
$F_{\nu}(t)=\frac{a^{2}}{D^{2}}\int d^{3}r\,n_{d}(r)\,\pi B_{\nu}(T_{d}(r,\theta,t))Q_{\nu}.$ (4)
@@
Intuitively, each point on the isodelay surface contributes a flux that can be thought of as the solution to the radiative transfer equation for a blackbody source function in the optically thin limit, where $F_{\nu}(t,\theta,t)\propto B_{\nu}(T_{d}(r,\theta,t))\,\tau_{\nu}$ and depends primarily on the dust temperature. By summing along the isodelay surface, we find
$F_{\nu}(t)=\frac{2\pi\sigma_{\nu}}{D^{2}}\int_{r_{\mathrm{sub}}}^{r_{\mathrm{out}}}dr\,r^{2}n_{d}(r)B_{\nu}\left(T_{d}\left(r,\theta=\theta_{0},t\right)\right),$ (5)
where $\sigma_{\nu}=\pi a^{2}Q_{\nu}$ is the dust cross-section and $\theta_{0}=\arccos\left(1-\frac{ct}{r}\right)$.
@@
Figure 5: Optically thin dust model for a $\delta$-function flare at time $t=0$. Each panel shows how the model changes for a different free parameter. The free parameters are the time since the flare ($\Delta t$), luminosity of the flare ($L$), density at the sublimation radius ($n_{\mathrm{sub}}$), and slope of the radial density profile ($\gamma$). The black model is a base model that is the same in each panel. The parameters for the model are: $\Delta t=5$ yr, $L=10^{44}$ erg s-1, $n_{\mathrm{sub}}=10^{-10}$ cm-3, and $\gamma=1.5$. All models assume a single grain size of $a=0.1$ µm  and an MRN mixture of silicates and graphite dust.
@@
Figure 5 shows how our model depends on each of the free parameters, which are the time since the flare ($\Delta t$), luminosity of the flare ($L$), density at the sublimation radius ($n_{\mathrm{sub}}$), and slope of the radial density profile ($\gamma$). In all four panels, the black model has $\Delta t=5$ years, $L=10^{44}$ erg s-1, $n_{\mathrm{sub}}=10^{-10}$ cm-3, $\gamma=1.5$, $a=0.1$ µm. Qualitatively, this model matches the late-time TDE spectra from JWST MIRI, with strong silicate emission features and comparable continuum shapes. The spectrum exhibits a clear reddening with time (top left), which is the result of seeing cooler, more distant dust at later times. Both the input luminosity of the flare and $\gamma$ affect the total observed flux and the spectral slope. Increasing the luminosity will lead to higher dust temperatures and therefore a bluer spectrum (top right), while increasing $\gamma$ leads to a more centrally concentrated density profile and a slightly bluer spectrum (bottom right), although this effect is more subtle. As long as the dust remains optically thin in the MIR, the density at the sublimation radius (i.e. the normalization) does not change the spectral slope, just the overall normalization (bottom left). The same is true for the dust grain size, $a$, which we do not show in this figure for brevity. This figure highlights that the parameter space is already quite complex, but the model can provide a qualitatively good explanation of the JWST MIRI spectra.
@@
Initial fits to the JWST MIRI data in the long wavelength range ($\lambda>8$ µm) reveal that our optically thin, isodelay surface model provides a good fit in this regime. The right panel of Figure 6 shows an example model for WTP14adbjsh, the closest and oldest TDE in our sample. The green, distant dust model, which arises from the green isodelay surface shown in the left panel, provides a good fit to the long wavelength spectrum ($\lambda>8$ µm). In this single $\delta$-function flare model, the short wavelength flux is systematically under-predicted, likely as a result of not including the long-lived nature of the TDE flare. A long-lived flare will show emission from dust at a range of isodelay surfaces, including emission from hotter dust closer to the SMBH. However, modeling the emission from close to the SMBH requires taking into account two important factors: (1) the dust close to the SMBH is likely dominated by graphite because graphite sublimates at a higher temperature than silicates (Barvainis, 1987), and (2) smaller dust grains are easier to destroy than larger dust grains (Schartmann et al., 2008), leading to larger dust grains dominating the MIR spectrum in close proximity to the SMBH. While a self-consistent prescription of these two effects is beyond the scope of this work, we show that the inclusion of pure graphite dust at 0.1 pc from the SMBH (i.e. near the sublimation radius; $\Delta t\approx 0.3$ years) provides an adequate description of the short wavelength spectrum. The graphite spectrum shown in purple in Figure 6 assumes the dust grains are larger with $a=0.3$ µm  and that the flare luminosity has dropped by a factor of 25 compared to the green model. The dust grain size and current luminosity are degenerate, but we will explore this in future work, since a fully self-consistent long-lived flare model will be necessary to probe this degeneracy accurately. While we only show models for WTP14adbjsh (the closest), these results are similar for all sources in our sample, as Figure 1 suggest that all four sources show a second hot component at short wavelengths. Future work will address this second component with more complex, self-consistent dust model.
Figure 6: Left: A 2D slice of the isodelay surface as a function of time with 0.3 and 9.3 years shown in purple and green, respectively. In this slice, the observer is at $y=+\infty$. The black circle roughly corresponds to the dust sublimation radius for a flare with $L=10^{44}$ erg s-1. With increasing time, we are seeing dust further from the SMBH. Right: Representative model to describe the continuum shape of WTP14adbjsh. The green line shows the distant dust model at a time $\Delta t=9.3$ years since the flare, with $L=10^{44}$ erg s-1, $n_{\mathrm{sub}}=6\times 10^{-10}$ cm-3, and $\gamma=1.2$. This fits the long wavelength spectrum well, but underestimates the short wavelength spectrum. The purple line shows that graphite grains near the SMBH can account for this short wavelength excess. In this purple model, we keep the density profile parameters ($n_{\mathrm{sub}}$ and $\gamma$) fixed to the values from the green model, assume the dust is purely graphite with a grain size of $a=0.3$ µm, and find that a luminosity that is 25$\times$ lower than initial flare is sufficient to reproduce the short wavelength spectrum. In reality, there will be emission from dust throughout the circumnuclear environment, so we stress that this is not a self-consistent model or direct fit to the data.

### v2

@@ Abstract
Mid-infrared (MIR) emission from tidal disruption events (TDEs) is a powerful probe of the circumnuclear environment around dormant supermassive black holes. This emission arises from the reprocessing of intrinsic emission into thermal MIR emission by circumnuclear dust. While the majority of optical- and X-ray-selected TDEs show only weak dust echoes consistent with primarily unobscured sight lines, there have been growing efforts aimed at finding TDEs in obscured environments using MIR selection methods. In this work, we present the first JWST observations of 4 MIR-selected TDEs with JWST Mid-Infrared Instrument (MIRI) Medium-Resolution Spectrometer (MRS). Two of these sources show flares in other wavelength bands (one in optical, one in X-ray), while the other two are MIR-only transients. None of these TDEs showed pre-outburst nuclear activity, but all of the MIRI/MRS observations reveal emission lines associated with highly ionized gas in the nucleus, implying ionization from TDE accretion. Additionally, all four sources show silicate emission features around 10 and 18 µm  that are much stronger than the features seen in active galactic nuclei (AGN). We suggest that the emitting dust is optically thin to its own emission and show that the MIR spectrum is consistent with emission from optically thin dust in the nucleus. All four sources show an excess at short wavelengths ($\lambda<8$ µm), which could arise from a late-time plateau in the intrinsic flare, akin to what is seen in late-time UV observations of unobscured TDEs, although self-consistent dust modeling is required to fully assess the strength of this late-time plateau.
@@ 1 Introduction
In the past 50 years, tidal disruption events (TDEs) have gone from a theoretical concept (Hills, 1975; Rees, 1988; Evans & Kochanek, 1989; Cannizzo et al., 1990) to an observable phenomenon (for a recent review, see Gezari, 2021), becoming one of the leading methods to study the dominant fraction of dormant supermassive black holes (SMBHs). TDEs occur when a star on a nearly radial orbit gets so close to an SMBH that the tidal forces overcome the star’s self-gravity. A fraction of the star’s mass falls back onto the SMBH, producing a flare of accretion that can be observed across the electromagnetic spectrum. In the late 1990s, the first TDEs were detected with transient soft, thermal X-ray emission, indicative of the formation of an accretion flow (e.g. Bade et al., 1996; Komossa & Bade, 1999; Komossa & Greiner, 1999; Donley et al., 2002; Esquej et al., 2007). A few more TDEs were discovered in the early 2000s with time-domain searches in the ultraviolet (UV) band using GALEX (e.g. Gezari et al., 2006, 2008, 2009). The discovery space is now dominated by time-domain optical surveys, which have increased the discovery rate by a factor of 10 already (van Velzen et al., 2011; Gezari et al., 2012; van Velzen et al., 2021b; Hammerstein et al., 2023; Yao et al., 2023) and are expected to increase the rate again by at least a factor of 10 with next generation surveys like the Legacy Survey for Space and Time (LSST) at the Vera Rubin Observatory (Bricman & Gomboc, 2020).
These canonical detection methods tend to find TDEs in relatively gas- and dust-poor environments, as evidenced by the low covering factors measured with TDE dust echoes in the mid-infrared (MIR; van Velzen et al., 2016; Jiang et al., 2021a). However, in gas- and dust-rich environments, the nuclear dust may have a larger covering factor, making it more likely that the dust covers the sight line to the nucleus. Dust around accreting SMBHs reprocesses the higher energy optical/UV/X-ray photons into thermal MIR emission (Lu et al., 2016; Jiang et al., 2016; van Velzen et al., 2016, 2021a), and thus, the MIR may be one of the few ways to detect TDEs in gas- and dust-rich environments. Initial searches for MIR flares in ultraluminous infrared galaxies (U/LIRGs) revealed a handful of TDE-like events (e.g. Mattila et al., 2018; Kool et al., 2020), suggesting an elevated rate of TDEs in obscured environments (Reynolds et al., 2022), although many of these showed some level of activity from an active galactic nucleus (AGN) prior to the MIR flare. Other efforts have involved searching for MIR flares in integrated photometry from the NEOWISE mission (Mainzer et al., 2014), using spectroscopically confirmed galaxies from SDSS (Jiang et al., 2021b). This approach revealed many potential TDE candidates, but was also agnostic to past AGN activity, thereby introducing many AGN contaminants into the sample.
@@ 2 Observations & Data Reduction
Figure 2: Left: Host galaxies of the four TDEs in our sample, adapted from Figure 9 in Masterson et al. (2024). Right: Multi-wavelength light curves of the four TDEs in our sample, adapted from Figure 1 in Masterson et al. (2024), including the most recent WISE data up to 2024. The left $y$-axis denotes the flux density in the IR and optical bands. The right $y$-axis denotes the resulting luminosity, corresponding to $\nu L_{\nu}$ at the central frequency of the $W2$ band for the IR band, the 0.2-2.3 keV luminosity for eROSITA, and the 0.2-2 keV luminosity for XMM-Slew. Red pentagons and orange diamonds denote the ATLAS $c$- and $o$-band data, respectively. The data are binned to 30 days, except for WTP18aampwj (the optical transient), which is binned to 10 day cadence to highlight the optical transient. Green $\times$’s and yellow hexagons show the X-ray data for WTP17aamzew (the X-ray transient), from eROSITA and XMM-Slew, respectively. The vertical black dashed line denotes the time of the JWST observation of each source.
@@ 4.2 Time-Dependent Optically Thin Dust Modeling
where $\overline{Q}_{\nu}$ is the average absorption efficiency integrated over the source spectrum. In the UV and optical band, $\overline{Q}_{\nu}\approx 1$, while $\overline{Q}_{\nu}$ is roughly 2-3 orders of magnitude lower in the IR. For relatively conservative choices of the density profile (e.g. constant density profile with $n_{\mathrm{d}}=10^{-9}$ cm-3), the dust is indeed optically thin to its own emission for the size scales we are investigating here ($r\sim$ few pc), regardless of the order of magnitude uncertainties on the gas density in galactic centers.
Motivated by the silicate emission features and this order of magnitude estimate, we modeled the JWST MIRI spectra with optically thin dust, neglecting the geometric effects and scattering, but accounting for the time-dependent effects from TDEs. For this initial investigation, we adopt dust composed of 53% silicate and 47% graphite grains (i.e. an MRN mixture, typical for the Milky Way’s ISM; Mathis et al., 1977) with a single size of $a=0.1$ µm. This choice of an MRN dust mixture is motivated by the presence of both silicates and PAHs (i.e. carbon-rich dust) in the MIR spectra. We use the astronomical silicate and graphite optical properties from Draine & Lee (1984) and Laor & Draine (1993) [^2] See https://www.astro.princeton.edu/~draine/dust/dust.diel.html. To model the underlying flare, we assume that the bolometric luminosity follows the standard optical TDE light curve, and thus, we adopt the post-peak form from Equation (3) in van Velzen et al. (2021b) with $p=5/3$ and an average value of $\log_{10}(t_{0})=1.75$ from optical flares (van Velzen et al., 2021b). We neglect the rising part of the light curve for this initial analysis and simply rescale the peak luminosity of the light curve with a free parameter, $L$. Motivated by recent observations of late-time plateaus in optical and UV data (e.g. Mummery et al., 2024), we assume that the luminosity remains constant once it reaches a luminosity of $L_{\mathrm{plat}}$.
To compute the response of the dust to this flare, we set up a spherically symmetric shell of dust from $r_{\mathrm{sub}}$ to $r_{\mathrm{out}}$, approximating the sublimation radius with Equation (1) of van Velzen et al. (2021a) with the initial flare luminosity and setting $r_{\mathrm{out}}=10$ pc. At each point in this shell, we numerically compute the expected temperature of the dust by solving the energy balance equation, given by
@@
where $L$ is the bolometric luminosity emitted by the source at the retarded time given by $t_{r}=t_{\mathrm{obs}}-\frac{r}{c}\left(1-\cos\theta\right)$, $\tau$ is the integrated optical depth given by Equation (2), $Q_{\nu}$ is the absorption efficiency, and $T_{d}(r,\theta,t_{r})$ is the dust temperature. At the locations where $t_{r}<0$, the dust has yet to respond to the flare, and hence there is no transient emission from this dust. This prescription considers only the radiative cooling of the dust, ignoring cooling from grain sublimation, which is only dominant at significantly higher temperatures than probed here ($T\gtrsim 2800$ K; e.g. Waxman & Draine 2000; Lu et al. 2016). Given the unknown source spectrum and the high efficiency of dust absorption at optical and UV wavelengths, we assume $\overline{Q}_{\nu}=1$. We adopt a power-law density profile given by $n(r)=n_{\mathrm{sub}}(r/r_{\mathrm{sub}})^{-\gamma}$, where the subscript “sub” denotes values at the sublimation radius.
At each point in our shell, we compute the volume emissivity, given by
$J_{\nu}=\pi a^{2}Q_{\nu}\,n_{d}(r)\,B_{\nu}\left(T_{d}(r,\theta,t)\right),$ (4)
which has units of erg s-1 cm-3 sr-1 Hz-1. The total dust emission is then found by integrating the emissivity over the entire shell volume, which yields
[…]
@@
This is akin to the approach presented in Lu et al. (2016), but with two modifications: (1) we preserve the spectral information, whereas Lu et al. (2016) use Planck-averaged emissivities, and (2) we do not account for the additional attenuation by the outer dust shells (i.e. Equations (18-21) in Lu et al. 2016). Intuitively, each point contributes a flux that can be thought of as the solution to the radiative transfer equation for a blackbody source function in the optically thin limit, where $F_{\nu}(t,\theta,t)\propto B_{\nu}(T_{d}(r,\theta,t))\,\tau_{\nu}$ and depends primarily on the dust temperature.
@@
Figure 6: Optically thin dust model for the late-time MIR spectrum from a TDE-like flare. Each panel shows how the model changes for a different free parameter. The free parameters are the time since the flare ($\Delta t$), luminosity of the flare ($L$), density at the sublimation radius ($n_{\mathrm{sub}}$), and slope of the radial density profile ($\gamma$). The black model is a base model that is the same in each panel. The parameters for the model are: $\Delta t=5$ yr, $L=5\times 10^{44}$ erg s-1, $n_{\mathrm{sub}}=10^{-10}$ cm-3, and $\gamma=1$. All models assume a single grain size of $a=0.1$ µm, an MRN mixture of silicates and graphite dust, and a plateau luminosity of $L_{\mathrm{plat}}=10^{42}$ erg s-1.
@@
Figure 6 shows how our model depends on each of the free parameters, which are the time since the flare ($\Delta t$), luminosity of the flare ($L$), density at the sublimation radius ($n_{\mathrm{sub}}$), and slope of the radial density profile ($\gamma$). In all four panels, the black model has $\Delta t=5$ years, $L=5\times 10^{44}$ erg s-1, $n_{\mathrm{sub}}=10^{-10}$ cm-3, $\gamma=1$, $a=0.1$ µm, and $L_{\mathrm{plat}}=10^{42}$ erg s-1. Qualitatively, this model matches the late-time TDE spectra from JWST MIRI, with strong silicate emission features and comparable continuum shapes. The spectrum exhibits a clear reddening with time (top left), which is the result of seeing cooler, more distant dust at later times. Both the input luminosity of the flare and $\gamma$ affect the total observed flux and the spectral slope. Increasing the luminosity will lead to higher dust temperatures and therefore a bluer spectrum (top right), while increasing $\gamma$ leads to a more centrally concentrated density profile and a slightly bluer spectrum (bottom right). As long as the dust remains optically thin in the MIR, the density at the sublimation radius (i.e. the normalization) does not appreciably change the spectral slope, just the overall normalization (bottom left). The same is true for the dust grain size, $a$, which we do not show in this figure for brevity. This figure highlights that the parameter space is already quite complex, but the model can provide a qualitatively good explanation of the JWST MIRI spectra.
@@
Initial comparisons to the JWST MIRI data reveal that our optically thin model provides a good fit in the long wavelength range ($\lambda>8$ µm). The bottom panel of Figure 7 shows an example model for WTP14adbjsh, the closest and oldest TDE in our sample. In this model, the short wavelength flux, which arises primarily from closer regions to the SMBH, is systematically under-predicted with the assumption of MRN dust at a single dust grain size. Modeling the emission from close to the SMBH requires taking into account two important factors: (1) the dust close to the SMBH is likely dominated by graphite because graphite sublimates at a higher temperature than silicates (Barvainis, 1987), and (2) smaller dust grains are easier to destroy than larger dust grains (Schartmann et al., 2008), leading to larger dust grains dominating the MIR spectrum in close proximity to the SMBH. While a self-consistent prescription of these two effects is beyond the scope of this work, we show that using larger pure graphite dust grains at late times (i.e. recently heated dust close to the SMBH) provides an adequate description of the short wavelength spectrum. We note that the grain size and plateau luminosity are degenerate, which we will explore in future work, since a fully self-consistent dust model will be necessary to probe this degeneracy accurately. While we only show models for WTP14adbjsh (the closest), these results are similar for all sources in our sample, as Figure 2 suggest that all four sources show a second hot component at short wavelengths.
Figure 7: Optically thin dust modeling of WTP14adbjsh (the closest). Top: Assumed underlying flare heating the dust, with $\Delta t=9.3$ years, $L=5.5\times 10^{44}$ erg s-1, $n_{\mathrm{sub}}=1.5\times 10^{-10}$ cm-3, and $\gamma=0$. The model plateaus when it reaches a luminosity of $L_{\mathrm{plat}}=2\times 10^{42}$ erg s-1. We denote the two phases in which we model the spectrum with MRN and graphite dust in green and purple, respectively. We assume that the graphite dust is responding to the last year of the flare to roughly mimic the dust composition evolution we expect in the innermost regions around the SMBH. Bottom: Resulting optically thin dust model from the above underlying accretion flare. The MRN dust (green curve) fits the long wavelength spectrum well, but underestimates the short wavelength spectrum. Larger graphite grains near the SMBH (purple curve) can account for this short wavelength excess. We stress that this is not a self-consistent model, nor direct fit to the data, and other parameter combinations could also produce a similarly good fit. We will fully explore this parameter space in a follow-up work with self-consistent dust physics incorporated.